OpenStudy (anonymous):

(f/g)(x), f(x)=square root of x, g(x)=2x+3. I set it up but, I don't know what to do next

OpenStudy (rsadhvika):

whats your setup ?

OpenStudy (anonymous):

i have (square root of x)/2x+3 then that times by x

OpenStudy (rsadhvika):

\[\large \left(\dfrac{f}{g}\right)(x) = \dfrac{f(x)}{g(x)} =\dfrac{\sqrt{x}}{2x+3}\]

OpenStudy (rsadhvika):

you're done. there is not to do next

OpenStudy (anonymous):

Oh okay, thank you

OpenStudy (anonymous):

@mayankdevnani yes it was